# FaaF: Facts as a Function for the evaluation of generated text

### v2

@@ Abstract
The demand for accurate and efficient verification of information in texts generated by large language models (LMs) is at an all-time high, but remains unresolved. Recent efforts have focused on extracting and verifying atomic facts from these texts via prompting LM evaluators. However, we demonstrate that this method of prompting is unreliable when faced with incomplete or inaccurate reference information. We introduce Facts as a Function (FaaF), a new approach to the fact verification task that leverages the function-calling capabilities of LMs. FaaF significantly enhances the ability of LMs to identify unsupported facts in texts, while also improving efficiency and significantly lowering costs compared to prompt-based methods. Additionally, we propose a framework for evaluating factual recall in Retrieval Augmented Generation (RAG) systems, which we employ to compare prompt-based and FaaF methods using various LMs under challenging conditions.
@@ 1 Introduction
In these applications, Retrieval Augmented Generation (RAG) is often used to integrate out-of-training knowledge to the LM (Lee et al., 2019). This is achieved by dynamically fetching and incorporating relevant data from a trusted reference source prior LM text generation, resulting in more informed, accurate, and contextually rich outputs.
Considering the importance of factual accuracy of generated text in this setting, a significant body of work has focused on automated ways for obtaining it. At a high level, two distinct schools of thought emerge - shaped by the intended use case, the available test datasets and required scale. First, authors who address factual precision —the truthfulness of each statement in a generated text— in both RAG and non-augmented LM generation scenarios (Chen et al., 2022; Zhang et al., 2023; Gao et al., 2023; Lee et al., 2023; Min et al., 2023; Azaria and Mitchell, 2023; Yuan et al., 2021; Fu et al., 2023; Wang et al., 2023; Kadavath et al., 2022; Manakul et al., 2023). Given that a generated text may include both accurate and inaccurate statements, a common approach is to initially extract a number of fact statements from the generated text and verify them individually by prompting an LM evaluator model with appropriate references. Secondly, other studies (Cuconasu et al., 2024; Liu et al., 2023; Kandpal et al., 2023; Mallen et al., 2023) attempt to evaluate LMs and RAG systems in terms of exact matching any one of a set of predefined correct answers in the generated text.
There is an apparent lack of work on the practical measurement of factual recall—the extent to which all the information required to sufficiently answer a question is included in the generated text. Factual recall is a key performance metric, particularly in the RAG scenario, because it directly captures the performance of retrieval and generation simultaneously and closely reflects the central purpose of the system. Factual recall is at least as important as factual precision for RAG systems, where a response can be factually precise but given the wrong context may be irrelevant to the posed question.
Although the exact match method can be viewed as a fast and efficient form of binary recall score (where the existence of a single accepted answer in the generated text signals success), it faces serious challenges and limitations in the verification task.
As recognised by Cuconasu et al. (2024), exact matching of ground truth answers in the generated text is prone to false negatives since the ground truth information might be present but phrased differently. Additionally, this approach cannot be applied when the grounding information is longer than a few words since the chances of exact match become very slim. Lastly, being binary renders it a relatively poor metric considering that there may be plural grounding information which need to be included in the generated text for a complete answer.
@@
To address the gap above, we make the following contributions:
@@
    We introduce Facts as a Function (FaaF), a new fact verification formulation which outperforms fact verification via prompting and reduces the required number of LM calls and completion tokens by more than 5 times.
@@
    We propose an end-to-end factual recall evaluation framework which is tailored to RAG systems. It can be used to (i) create a test dataset and (ii) perform automated factual recall evaluation given a RAG system (or generally a language model).
@@
    We probe into the performance of fact verification formulations in conditions of highly incomplete or inaccurate generated text. To achieve that, we augment WikiEval [^1] https://huggingface.co/datasets/
    explodinggradients/WikiEval
    (Es et al., 2023) with ground truth facts and human annotation of their validity. WikiEval features question/answer pairs with answers of variable factual quality which enable simulating deficient RAG responses. We find that prompt-based fact verification faces serious challenges in identifying unsupported facts in the presence of inaccurate or incomplete generated text.
  4. 4.
@@ 2 Related Work
Recently, authors introduced RAGAS (Es et al., 2023), an evaluation framework for RAG systems which is measuring the performance of retrieval and generation without the requirement of ground truth human annotation. Although the authors recognise and are motivated by the need for a practical an efficient RAG evaluation method, RAGAS does not capture the factual aspect of the evaluation, which is a key performance metric and the central focus of RAG’s intended use.
Min et al. (2023) use prompt variations and an aggregate non-parametric probability of the tokens in a fact statement to directly verify individual facts extracted from LM generated biographies. They compare their method with human evaluation and find low error rates when retrieving the ground truth for the evaluated fact.
@@
Azaria and Mitchell (2023) also find that fact verification by prompting insufficient and propose to train a classifier on the hidden layer activations of open source LMs to predict the truthfulness of a generated statements. However, current leading commercial models are lacking layer activation access and so require alternative methods.
Another approach in the same spirit is to look at the conditional probabilities of each generated token as an indicator of LM confidence and truthfulness of the generated text with the view that low LM confidence is a proxy for incorrect statements (Yuan et al., 2021).
@@
Lastly, recent work on the factual accuracy of LMs and RAG systems (Cuconasu et al., 2024; Liu et al., 2023; Kandpal et al., 2023; Mallen et al., 2023) took the approach of using the NaturalQuestions-Open (NQ-open) [^4] https://ai.google.com/research/NaturalQuestions dataset (Kwiatkowski et al., 2019) and calculate accuracy by judging whether any of the ground truth answers (NaturalQuestions annotations) appear in the generated text via exact matching. NQ-open is a large scale dataset which comprises historical Google search queries and their human-annotated answers sourced from Wikipedia. Even though NQ-open, is valuable for its extensive scope and domain-agnostic nature, fact-verification via exact matching faces serious challenges. As highlighted by Cuconasu et al. (2024), a key issue is accurately determining the truthfulness of answers, especially when dealing with phrases that have the same meaning but (slightly different format or wording) or different date formats. Thus, the necessity for a more advanced analysis of answer variations is recognised and left for future research.
@@ 3 Facts as a Function
Facts as a Function (FaaF) is a streamlined fact verification method using function calling for multi-fact assessment. Although in this paper we focus in examining FaaF in the scope of factual recall evaluation, the method itself can equivalently be used for factual precision and broader fact verification tasks.
Key idea 1: JSON and XML over prompt
We propose that by using the function calling ability of the LM, we enforce a more formal mode of token generation compared to natural language for three reasons. First, by leveraging the metadata of function arguments, type annotations and tailored instructions can effectively constrain the LM to the accepted modes of response. This provides an element of structured repetition of the instructions towards the LM (via the metadata attached to each function argument) and ultimately results in a more consistent guidance compared to instructions given once at the end or the beginning of a prompt. Secondly, with formalising the type annotation, we can avoid relying on exact matching to interpret the LM’s text response—which can prove detrimental as we demonstrate in the results of this paper. Type annotations can be combined with custom types to essentially convert a function argument into a classification result to a multiple-choice question. Building on the findings of Kadavath et al. (2022), who established that LMs show well calibrated probabilities when presented with multiple choice questions, we propose that using the function argument’s type annotations to convey the accepted LM responses is a step further in the same direction. Thirdly, due to the strict nature of code syntax compared to natural language, gradients during training are expected to be steeper and thus the LM learns to do a better job at adhering to the expected output and responding with lower stochasticity—which is particularly helpful in the fact verification use case.
Key idea 2: Generated text as a unit
As discussed above, a function definition can encapsulate a sufficient number of arguments to be used by the LM so that they capture all the facts statements which need to be verified. Therefore, we move away from the concept that each fact should be verified individually via a fact-specific prompt and we propose instantiating a function per LM-generated text which needs to be factually assessed. The function can be programatically generated in a manner such that it includes all the individual facts which need to be verified in a given piece of text. In other terms, function calling allows for the fact verification of a long-form text as a unit, with a single LM call. This approach results in a reduction of cost and time for fact-verification which is proportional to the number of facts which would otherwise needed to be assessed individually, as seen in Chen et al. (2022); Gao et al. (2023); Min et al. (2023); Lee et al. (2023).
Key idea 3: Outsourcing judgement from the LM to the function
@@
$M$ responds with the output $o_{x}$ which is of string type. Then, a parsing function $G_{M}$ which adheres to particular response schema of $M$ is used to parse the raw response $o_{x}$ and invoke $O$ by assigning values on its arguments, yielding $\acute{O}$. The values being assigned to each of $(f_{1},f_{2},\ldots,f_{n})$ being the verification result of the underlying fact statement. Upon invocation of $O$, the function arguments undergo type validation to ensure that its schema and type annotations are respected.
@@ 4 Assessment of fact verification formulations in the RAG setting
The ground truth answer is generated by providing the LM with the correct context from the respective Wikipedia page. The ungrounded answer is generated by not providing any context to the LM. Finally, the poor answer is generated by instructing the LM to give an incomplete answer to the given question. An example of each answer type can be seen in Figure 2. Generally, ungrounded answers contain false, incomplete and redundant information with respect to the ground truth answers. On the other hand, poor answers contain primarily incomplete information compared to ground truth answers i.e. no evidence for support or rejection for ground truth facts.
This dataset enables us to asses the impact of quality and completeness of different answer variants to the performance of fact-verification. In that way, we test closely the ability of different fact verification methods and LMs to identify unsupported facts when presented with (i) incorrect, (ii) indirectly relevant and (ii) incomplete information with some degree of distinction.
@@
For each set of fact statements which encapsulate a ground truth answer, we construct a facts-specific function object and a parsing function. Since the created function object contains all the input facts as arguments, we perform verification to the set of facts as a unit. An example of the JSON representation of a function object containing the first fact can be seen bellow:
{’properties’:
    ‘fact_0:{
    ’description’:"It
    is clear from the passage
    that Pope Benedict XVI became
    the head of the Catholic Church
    and sovereign of the Vatican
    City State on April 19, 2005.
    Respond by using one of the accepted
    Enum types.",
   Ψ’enum’: [’True’, ’False’],
   Ψ’type’: ‘string’
’required’:
Ψ[‘fact_0’,
 Ψ‘fact_1’
 Ψ …,
 Ψ‘fact_n’]
’title’: ’FactChecker’,
’type’: ‘object’}
Each function argument includes metadata which can be used to pass instructions, type annotations and the fact statement to be verified itself. In addition to the function object, we pass the following prompt:
@@
We use the latest commercially available models which support function calling gpt-4-turbo, gpt-3.5-turbo, claude-3-opus and claude-3-sonnet. We also examined the recently released (mistral-Large) [^5] https://mistral.ai/news/mistral-large/ LM but it was excluded from the results in this paper due to its high failure rate of over 80% in some cases in generating appropriately formatted responses, rendering its results non-contributory to the discussion. It is important to note that we did not allow models to retry in case of a failed parsing of their response or failed invocation of the function object due to formatting. FaaF introduces strict constraints on the expected LM response and by permitting only one attempt, we also assess the LM’s proficiency in formatting their response, as well as verifying factual accuracy.
In addition, we highlight the crucial role of system prompt to model performance. A change in the system prompt can impact significantly the fact verification accuracy of the LM. We kept the system prompts for the GPT models unchanged but modified Claude models’ system prompts to incorporate a 1-shot example of simple function calling. This adjustment is a result of following the official function-calling recommendations of the two model families at the time of writing this paper — GPT uses JSON, while Claude uses XML. As the pace of developments in this space is fast, we expect further changes and improvements in the optimal interface between LMs and tools/functions.
Lastly, a comprehensive assessment of open source models which support function calling is left for future study as there are none widely established available at this time.
@@
This paper’s metrics exclusively evaluate language models (LMs) based on their successfully formatted responses where applicable (FaaF). In doing so, we ensure that the comparison of the LMs’ fact verification ability is not influenced by their capacity to format responses correctly — which is discussed separately.
Error rate
We use Error Rate (ER) between the human fact-verification and the fact verification formulation as the main indicator of verification accuracy.
F1micro
We also use F1micro score (F1m) as defined in Min et al. (2023) to measure the successful identification of unsupported facts and probe further into the individual fact verification. It should be noted that F1 scores explicitly depend on the class ratio (T/F) via precision and recall. For that reason F1m scores should be compared across fact verification approaches in the same answer category (where the T/F ratio is preserved) in Table 2, and not across answer categories.
@@
gpt-3.5-turbo Prompt(T/F) 0/281 1.4 $*{}^{*}$ 0/281 27.4 76.5 16/281 55.3 56.9
FaaF(T/F) 0/281 1 $*{}^{*}$ 0/281 23.1 81 0/281 17.4 89.6
FaaF(T/F/N) 0/281 1 $*{}^{*}$ 0/281 25.9 78 0/281 18.5 88.8
FaaF(T/F)+citation 0/281 1 $*{}^{*}$ 26/281 28.2 75.3 71/281 15.7 90
FaaF(T/F/N)+citation 0/281 1.4 $*{}^{*}$ 26/281 31.3 71 31/281 23.2 85.5
claude-3-sonnet Prompt(T/F) 0/281 1 $*{}^{*}$ 0/281 42.7 58 2/281 77.7 26.4
FaaF(T/F) 0/281 1.7 $*{}^{*}$ 0/281 28.8 75.9 0/281 14.9 91.1
FaaF(T/F/N) 0/281 2 $*{}^{*}$ 0/281 27.7 76.9 0/281 14.2 91.6
FaaF(T/F)+citation 0/281 0.7 $*{}^{*}$ 0/281 30.6 73.9 0/281 14.5 91.4
FaaF(T/F/N)+citation 0/281 1 $*{}^{*}$ 0/281 27.4 76.7 0/281 14.9 91.3
gpt-4-turbo Prompt(T/F) 0/281 1.7 $*{}^{*}$ 1/281 24.2 80.1 9/281 44.1 68.4
FaaF(T/F) 0/281 1 $*{}^{*}$ 0/281 22.4 82 0/281 8.5 95.1
FaaF(T/F/N) 0/281 1.4 $*{}^{*}$ 0/281 17.7 86.3 0/281 6.7 96.2
FaaF(T/F)+citation 0/281 0.7 $*{}^{*}$ 6/281 15.6 88.2 15/281 7.5 95.8
FaaF(T/F/N)+citation 0/281 1 $*{}^{*}$ 6/281 16 87.7 0/281 9.2 94.8
claude-3-opus Prompt(T/F) 0/281 1.7 $*{}^{*}$ 0/281 42.3 58.8 2/281 76.3 28.7
FaaF(T/F) 0/281 3.2 $*{}^{*}$ 0/281 15.3 88.8 0/281 6.7 96.2
FaaF(T/F/N) 0/281 3.5 $*{}^{*}$ 0/281 14.5 89.3 0/281 4.9 97.2
FaaF(T/F)+citation 0/281 0.3 $*{}^{*}$ 0/281 24.1 80.8 0/281 5 97
FaaF(T/F/N)+citation 0/281 0.7 $*{}^{*}$ 0/281 20.9 83.6 0/281 7.8 95.6
Table 2: Results on Error Rate (ER) along with F1micro estimated by each fact formulation. F1micro measures accuracy and recall in predicting the false facts. $*{}^{*}$1micro is 0 in the Ground Truth Answer cases because there are no False facts in the ground truth evaluation in this category. T/F indicates True/False and T/F/N are True/False/Not clear from the given passage respectively. N/A indicates the Not answered i.e. the number of facts that the LM did not respond on their validity due to due to errors in its response format.

### v3

@@ Abstract
The demand for accurate and efficient verification of information in texts generated by large language models (LMs) is at an all-time high, but remains unresolved. Recent efforts have focused on extracting and verifying atomic facts from generated texts via prompting LM evaluators. However, we demonstrate that this method of prompting is unreliable when faced with incomplete or inaccurate reference information. We introduce Facts as a Function (FaaF), a new approach to the fact verification task that leverages the function-calling capabilities of LMs. FaaF significantly enhances the ability of LMs to identify unsupported facts in texts, while also improving efficiency and significantly lowering costs compared to prompt-based methods. Additionally, we propose a framework for evaluating factual recall in Retrieval Augmented Generation (RAG) systems, which we employ to compare prompt-based and FaaF methods using various LMs under challenging conditions.
@@ 1 Introduction
In these applications, Retrieval Augmented Generation (RAG) is often used to integrate out-of-training knowledge to the LM (Lewis et al., 2021).
Considering the importance of factual accuracy of generated text in this setting, a significant body of work has focused on automated ways for obtaining it. At a high level, two distinct schools of thought emerge - shaped by the intended use case, the available test datasets and required scale. First, authors who address factual precision —the truthfulness of each statement in a generated text— in both RAG and non-augmented LM generation scenarios (Chen et al., 2022; Zhang et al., 2023; Gao et al., 2023; Lee et al., 2023; Min et al., 2023; Azaria and Mitchell, 2023; Yuan et al., 2021; Fu et al., 2023; Wang et al., 2023; Kadavath et al., 2022; Manakul et al., 2023). Since a generated text may include both accurate and inaccurate statements, a common approach is to initially extract a number of fact statements from the generated text and verify them individually by prompting an LM evaluator model with appropriate references. Secondly, other studies (Cuconasu et al., 2024; Liu et al., 2023; Kandpal et al., 2023; Mallen et al., 2023) attempt to evaluate LMs and RAG systems in terms of exact matching any one of a set of predefined, correct answers in the generated text.
However, as recognised by Cuconasu et al. (2024), exact matching of ground truth answers in the generated text is prone to false negatives since the ground truth information might be present but phrased differently. This issue is exacerbated when the grounding information is longer than a few words making the conditional probability of an exact match very low. Secondly, merely the presence of specific words or phrases in the generated text is not sufficient to evaluate the truthfulness of more sophisticated statements.
Although the exact match method can be viewed as binary form of factual recall, where the existence of a single accepted answer in the generated text signals success, it faces serious limitations in the verification task. Considering the importance of factual recall —the extent to which all the information required to sufficiently answer a question is included in the generated text— in the RAG scenario, there is an apparent lack of work on its practical measurement. Factual recall directly captures the performance of retrieval and generation of a RAG simultaneously and closely reflects the central purpose of the system. In this use case, recall is more important than precision since a generated response can be factually precise but given the wrong context may be irrelevant to the posed question.
@@
To address the gaps above, we make the following contributions:
@@
    We introduce Facts as a Function (FaaF), a new fact verification formulation which results in significantly more accurate and efficient verification compared to current approaches relying on prompting.
@@
    We probe into the performance of fact verification formulations in conditions of highly incomplete or inaccurate generated text in a controlled manner. To achieve that, we augment WikiEval [^1] https://huggingface.co/datasets/explodinggradients/WikiEval(Es et al., 2023) which features question/answer pairs with answers of variable factual quality which enable simulating deficient RAG responses. We find that prompt-based fact verification faces serious challenges in identifying unsupported facts in the presence of inaccurate or incomplete generated text.
@@ 2 Related Work
Recently, Min et al. (2023) used prompt variations and an aggregate non-parametric probability of the tokens in a fact statement to directly verify individual facts extracted from LM generated biographies. They compare their method with human evaluation and find low error rates when retrieving the ground truth for the evaluated fact.
@@
Azaria and Mitchell (2023) also find that fact verification by prompting is insufficient and propose to train a classifier on the hidden layer activations of open source LMs to predict the truthfulness of a generated statements. However, current leading commercial models are lacking layer activation access and so require alternative methods.
Another approach in the same spirit is to look at the probabilities of each generated token as an indicator of LM confidence and truthfulness of the generated text with the view that low LM confidence is a proxy for incorrect statements (Yuan et al., 2021).
@@
Lastly, recent work on the factual accuracy of LMs and RAG systems (Cuconasu et al., 2024; Liu et al., 2023; Kandpal et al., 2023; Mallen et al., 2023) took the approach of using the NaturalQuestions-Open (NQ-open) [^4] https://ai.google.com/research/NaturalQuestions dataset (Kwiatkowski et al., 2019) and calculate accuracy by judging whether any of the ground truth answers (NaturalQuestions annotations) appear in the generated text via exact matching. NQ-open is a large scale dataset which comprises historical Google search queries and their human-annotated answers sourced from Wikipedia. Even though NQ-open is valuable for its extensive scope and domain-agnostic nature, fact-verification via exact matching faces serious challenges (Cuconasu et al., 2024) and a more advanced verification of answers is left for future research.
@@ 3 Facts as a Function
Facts as a Function (FaaF) is a streamlined fact verification method using function calling for multi-fact assessment.
Assuming a set of fact statements to be verified, we construct a facts-specific function object and a parsing function. Since the created function object contains all the input facts as arguments, we perform verification to the set of facts as a unit. An example of the JSON representation of a function object containing the first fact can be seen bellow:
{’properties’:
    ‘fact_0:{
    ’description’:"It
    is clear from the passage
    that Pope Benedict XVI became
    the head of the Catholic Church
    and sovereign of the Vatican
    City State on April 19, 2005.
    Respond by using one of the accepted
    Enum types.",
   Ψ’enum’: [’True’, ’False’],
   Ψ’type’: ‘string’
’required’:
Ψ[‘fact_0’,
 Ψ‘fact_1’
 Ψ …,
 Ψ‘fact_n’]
’title’: ’FactChecker’,
’type’: ‘object’}
Code over natural language
We propose that by using the function calling ability of the LM, we enforce a more formal mode of token generation compared to natural language. Function calling can be viewed as prompting the LM to generate code (the function arguments JSON) but this has important effect in the generation. First, due to the strict nature of code syntax compared to natural language, gradients during training are expected to be steeper which ultimately leads to better model calibration when generating code (e.g. arguments for a function). This manifests as better adhering to the expected output and responding with lower stochasticity.
By leveraging the metadata of function arguments, type annotations and tailored instructions we can constrain the LM to the accepted modes of response more effectively than prompting. Ultimately, we can avoid relying on exact matching to interpret the LM’s response—which can prove detrimental as we demonstrate in the results of this paper.
Additionally, type annotations can be combined with custom types to essentially convert a function argument into a classification result to a multiple-choice question. Building on the findings of Kadavath et al. (2022), who established that LMs show well calibrated probabilities when presented with multiple choice questions, we propose that using the function argument’s type annotations to convey the accepted LM responses is a step further in the same direction.
Generated text as a unit
A function definition can encapsulate all the facts statements which need to be verified for a piece of long-form text. Therefore, we move away from the concept that each fact should be verified individually via a fact-specific prompt and we propose instantiating a function per LM-generated text which needs to be factually assessed. This results in a reduction of cost and time for fact-verification which is proportional to the number of facts which would otherwise needed to be assessed individually, as seen in Chen et al. (2022); Gao et al. (2023); Min et al. (2023); Lee et al. (2023).
Outsourcing judgement from the LM to the function
@@
$M$ responds with the output $o_{x}$ which is passed to a parsing function $G_{M}$ which adheres to particular response schema of $M$ and invokes $O$ by assigning values on its arguments, yielding $\acute{O}$. The values being assigned to each of $(f_{1},f_{2},\ldots,f_{n})$ being the verification result of the underlying fact statement. The generated function arguments undergo type validation and modification, if required, to ensure that the schema and type annotations are respected.
@@ 4 Assessment of fact verification formulations in the RAG setting
The ground truth answer is generated by providing the LM with the correct context from the respective Wikipedia page. The ungrounded answer is generated by not providing any context to the LM. Finally, the poor answer is generated by instructing the LM to give an incomplete answer to the given question. An example of each answer type can be seen in Figure 2. Generally, ungrounded answers contain false, incomplete and redundant information with respect to the ground truth answers. On the other hand, poor answers contain primarily incomplete information or lack information altogether, compared to ground truth answers i.e. no evidence for support or rejection for ground truth facts.
This dataset enables us to assess the impact of quality and completeness of different answer variants to the performance of fact-verification. In that way, we test closely the ability of different fact verification methods and LMs to identify unsupported facts when presented with (i) incorrect, (ii) indirectly relevant and (ii) incomplete information with some degree of distinction.
@@
For each set of fact statements which encapsulate a ground truth answer, we construct a facts-specific function object as seen in section 3. Each function argument includes metadata which can be used to pass instructions, type annotations and the fact statement to be verified itself. In addition to the function object, we pass the following prompt:
@@
We use established commercially available models which support function calling gpt-4-turbo, gpt-3.5-turbo, claude-3-opus and claude-3-sonnet. We also examined mistral-Large but it was excluded from the results in this paper due to its high failure rate of over 80% in some cases in generating appropriately formatted responses, rendering its results non-contributory to the discussion. It is important to note that we did not allow models to retry in case of a failed invocation of the function object due to formatting. FaaF introduces strict constraints on the expected LM response and by permitting only one attempt, we also assess the LM’s proficiency in formatting their response, as well as verifying factual accuracy.
We kept the system prompts for the GPT models unchanged but modified Claude LMs system prompts to incorporate a 1-shot example of simple function calling. This adjustment is a result of following the official function-calling recommendations of the two model families at the time of writing this paper.
@@
The capacity of the LM to return a correctly formatted response for function calling is distinct to their ability for accurate fact verification. The verification accuracy metrics are calculated considering only the correctly formatted LM responses (first attempt only). In doing so, we ensure that the comparison of the LMs’ fact verification ability is not influenced by their capacity to format responses correctly — which is discussed separately. We use Error Rate (ER) between the human fact-verification and the fact verification formulation as the main indicator of verification accuracy. We also use F1micro score (F1m) as defined in Min et al. (2023) to measure the successful identification of unsupported facts and probe further into the individual fact verification. It should be noted that F1 scores explicitly depend on the class ratio (T/F) via precision and recall. For that reason F1m scores should be compared across fact verification approaches in the same answer category (where the T/F ratio is preserved) in Table 2, and not across answer categories.
@@
gpt-3.5-turbo Prompt(T/F) 0/281 1.4 0 0/281 27.4 76.5 16/281 55.3 56.9
FaaF(T/F) 0/281 1 0 0/281 23.1 81 0/281 17.4 89.6
FaaF(T/F/N) 0/281 1 0 0/281 25.9 78 0/281 18.5 88.8
FaaF(T/F)+citation 0/281 1 0 26/281 28.2 75.3 71/281 15.7 90
FaaF(T/F/N)+citation 0/281 1.4 0 26/281 31.3 71 31/281 23.2 85.5
claude-3-sonnet Prompt(T/F) 0/281 1 0 0/281 42.7 58 2/281 77.7 26.4
FaaF(T/F) 0/281 1.7 0 0/281 28.8 75.9 0/281 14.9 91.1
FaaF(T/F/N) 0/281 2 0 0/281 27.7 76.9 0/281 14.2 91.6
FaaF(T/F)+citation 0/281 0.7 0 0/281 30.6 73.9 0/281 14.5 91.4
FaaF(T/F/N)+citation 0/281 1 0 0/281 27.4 76.7 0/281 14.9 91.3
gpt-4-turbo Prompt(T/F) 0/281 1.7 0 1/281 24.2 80.1 9/281 44.1 68.4
FaaF(T/F) 0/281 1 0 0/281 22.4 82 0/281 8.5 95.1
FaaF(T/F/N) 0/281 1.4 0 0/281 17.7 86.3 0/281 6.7 96.2
FaaF(T/F)+citation 0/281 0.7 0 6/281 15.6 88.2 15/281 7.5 95.8
FaaF(T/F/N)+citation 0/281 1 0 6/281 16 87.7 0/281 9.2 94.8
claude-3-opus Prompt(T/F) 0/281 1.7 0 0/281 42.3 58.8 2/281 76.3 28.7
FaaF(T/F) 0/281 3.2 0 0/281 15.3 88.8 0/281 6.7 96.2
FaaF(T/F/N) 0/281 3.5 0 0/281 14.5 89.3 0/281 4.9 97.2
FaaF(T/F)+citation 0/281 0.3 0 0/281 24.1 80.8 0/281 5 97
FaaF(T/F/N)+citation 0/281 0.7 0 0/281 20.9 83.6 0/281 7.8 95.6
Table 2: Results on Error Rate (ER) along with F1micro estimated by each fact formulation. F1micro measures accuracy and recall in predicting the false facts. Bold case indicates best performance. F1micro is 0 in the Ground Truth Answer cases because there are no False facts in the ground truth evaluation in this category. T/F indicates True/False and T/F/N are True/False/Not clear from the given passage respectively. N/A indicates the Not answered i.e. the number of facts that the LM did not respond on their validity due to errors in its response format.